# Periodic orbits of non-degenerate lacunary contact forms on prequantization bundles

### v2

@@ 4.2. Proof of the theorem
#$\#\{1\leq i\leq r;\ {\mu}(\gamma_{i}^{k_{i}})<d-1\}+\#\{1\leq i\leq r;\ {\mu}(\gamma_{i}^{k_{i}})>d+1\}=r_{B}-\dim{\mathrm{H}}_{n}(B;\mathbb{Q}).$
@@
#$\#\{1\leq i\leq r;\ {\mu}(\gamma_{i}^{k_{i}})<d-1\text{ or }{\mu}(\gamma_{i}^{k_{i}})>d+1\text{ or }{\mu}(\gamma_{i}^{k_{i}})=d\}=r.$
@@
Let us first prove that $\alpha$ has precisely $r_{B}$ contractible closed orbits. First of all, note that ${\widebar{M}}$ does not need to have a symplectic filling. Thus, we will argue as in the proof of Theorem 1.6 under the assumption (NF), but for this we have to show that $\alpha$ is index-admissible. Let ${\hat{\alpha}}$ be the lift of $\alpha$ to $M$.
@@
Let ${\mathscr{P}}^{0}({\hat{\alpha}})$ and ${\mathscr{P}}^{0}(\alpha)$ be the set of (not necessarily simple) contractible closed orbits of ${\hat{\alpha}}$ and $\alpha$ respectively. Then there exists a map $\psi:{\mathscr{P}}^{0}({\hat{\alpha}})\to{\mathscr{P}}^{0}(\alpha)$ such that ${\mu}({\hat{\gamma}})={\mu}(\psi({\hat{\gamma}}))$. Moreover, this map has an inverse on the right $\rho:{\mathscr{P}}^{0}(\alpha)\to{\mathscr{P}}^{0}({\hat{\alpha}})$.
@@
Let ${\hat{\gamma}}$ be a contractible closed orbit of ${\hat{\alpha}}$ and $\tau:M\to{\widebar{M}}$ be the quotient projection. Define $\psi({\hat{\gamma}})=\tau\circ{\hat{\gamma}}$. Clearly, this is a contractible closed orbit of $\alpha$ with the same index. To construct $\rho$, let $\gamma$ be a contractible closed orbit of $\alpha$ and choose a point $x_{0}\in\tau^{-1}(\gamma(0))$. Since $\tau:M\to{\widebar{M}}$ is a finite covering and $\gamma$ is contractible, we have that $\gamma$ admits a closed lift ${\hat{\gamma}}$ such that ${\hat{\gamma}}(0)=x_{0}$. This lift is also contractible because given a capping disk $f$ of $\gamma$ (i.e. a continuous map $f:D^{2}\to{\widebar{M}}$ such that $f|_{\partial D^{2}}=\gamma$) we have that $f$ admits a lift $\hat{f}$ to $M$ such that $\hat{f}|_{\partial D^{2}}={\hat{\gamma}}$. Define $\rho(\gamma)={\hat{\gamma}}$. Then $\psi(\rho(\gamma))=\psi({\hat{\gamma}})=\gamma$ as desired. ∎
It follows from the first statement of the previous lemma that ${\hat{\alpha}}$ is lacunary. Since $M$ has a filling satisfying the assumption (F) of Theorem 1.6, we can consider the positive equivariant symplectic homology of $M$. As explained in Section 2.1, this is the homology of the complex ${\mathrm{CC}}^{0}_{*}({\hat{\alpha}})$ generated by the contractible good orbits of ${\hat{\alpha}}$. Since ${\hat{\alpha}}$ is lacunary, every periodic orbit is good and the differential of ${\mathrm{CC}}^{0}_{*}({\hat{\alpha}})$ vanishes. Therefore,
@@
for every contractible closed orbit ${\hat{\gamma}}$ of ${\hat{\alpha}}$. But, by (2.1),
@@
which implies that ${k_{\text{m}in}}=2c_{B}-n$. We claim that ${k_{\text{m}in}}>3-n$. Indeed, since $c_{B}>n/2$, ${k_{\text{m}in}}>n-n\geq 3-n$ if $n\geq 3$. If $n=2$, $c_{B}>1\implies 2c_{B}>2\implies 2c_{B}\geq 4\implies{k_{\text{m}in}}\geq 2>3-n$. If $n=1$, then $B=S^{2}$ which implies that $c_{B}=2\implies{k_{\text{m}in}}=3>3-n$. Hence, ${\hat{\alpha}}$ is index-admissible and so is $\alpha$, since the map $\psi$ in Lemma 5.1 is surjective.
@@
${\mathrm{HC}}^{0}_{\ast}({\widebar{M}})\cong\bigoplus_{m\in\mathbb{N}}{\mathrm{H}}_{\ast-2mc_{B}+n}(B;{\Lambda}).$ (5.1)
@@
${\mathrm{CC}}^{0}_{*}(\eta)\cong\bigoplus_{m\in\mathbb{N}}{\mathrm{H}}_{\ast-2mc_{B}+n}(B;{\Lambda})$ (5.2)
@@
Let ${\widebar{\xi}}=\tau_{*}\xi$ be the contact structure on ${\widebar{M}}$, where $\tau:M\to{\widebar{M}}$, as before, is the quotient projection. To define the equivariant symplectic homology of contractible orbits ${\mathrm{HC}}^{0}_{*}({\widebar{M}})$ using the symplectization of ${\widebar{M}}$ (with an integral grading) it is enough to have that $c_{1}({\widebar{\xi}})|_{\pi_{2}({\widebar{M}})}=0$: we do not need that $c_{1}({\widebar{\xi}})|_{H_{2}({\widebar{M}};\mathbb{Q})}=0$ as in assumption (NF). Indeed, this assumption is needed in Theorem 1.6 in order to compute ${\mathrm{HC}}^{0}_{*}({\widebar{M}})$ in item (b) of Theorem 2.1. Here, we do not need this since we compute ${\mathrm{HC}}^{0}_{*}({\widebar{M}})$ directly from the construction of $\eta$. The condition $c_{1}({\widebar{\xi}})|_{\pi_{2}({\widebar{M}})}=0$ holds because, by the assumption that $(M,\xi)$ satisfies (F), we have that $c_{1}(\xi)|_{\pi_{2}(M)}=0$ which is equivalent to the condition $c_{1}({\widebar{\xi}})|_{\pi_{2}({\widebar{M}})}=0$ since the induced map $\tau_{\#}:\pi_{2}(M)\to\pi_{2}({\widebar{M}})$ is an isomorphism.
@@
Since $M$ is closed, this action has finitely many isotropy subgroups $H_{1},\dots,H_{m}$. At most one of these subgroups has dimension $d$, say $H_{m}$. Since the codimension of all other stabilizers $H_{1},\dots,H_{m-1}$ is positive, we can find a compact 1-parameter subgroup of $T^{d}$ which is not contained in any $H_{i}$ with $1\leq i<m$. This is the desired subgroup. ∎
@@
where $R_{\beta}$ is the Reeb flow of $\beta$ (that generates the circle action of the prequantization $M$ whose orbits are the fibers) and $X_{i}^{h}$ is the horizontal lift of $X_{i}$; see [6, Lemma 3.4]. The Reeb flow of $R_{i}$ generates a circle action [6, Proposition 3.6] and commutes with the Reeb flow of $R_{\beta}$ [6, Lemma 3.8]. We also have that $[R_{i},R_{j}]=0$ for every $i$ and $j$. As a matter of fact,
@@
where $[{\hat{H}}_{i}R_{\beta},{\hat{H}}_{j}R_{\beta}]=0$ because ${\hat{H}}_{i}$ and ${\hat{H}}_{j}$ are invariant by the flow of $R_{\beta}$, $[{\hat{H}}_{i}R_{\beta},X_{j}^{h}]=0$ (and similarly $[X_{i}^{h},{\hat{H}}_{j}R_{\beta}]=0$) since $[R_{\beta},X_{j}^{h}]=0$ (see [6, Lemma 3.8]) and $X_{j}^{h}({\hat{H}}_{i})=X_{j}(H_{i})=\{H_{j},H_{i}\}=0$, $[X_{i},X_{j}]^{h}=0$ is the horizontal lift of $[X_{i},X_{j}]$, and $\omega(X_{i},X_{j})=0$ because $X_{i}$ and $X_{j}$ commute.
@@
Now, let $\epsilon>0$ be an irrational number and define $\eta_{1}^{\epsilon}=\beta/({\hat{H}}_{i}+\epsilon)$ so that the Reeb vector field of $\eta_{1}^{\epsilon}$ is $R_{1}^{\epsilon}=R_{1}+\epsilon R_{d+1}$. By the previous discussion, we have that the $T^{d+1}$-action preserves $\eta^{\epsilon}_{1}$ as well: given $g\in T^{d+1}$ we have that $g^{*}\eta^{\epsilon}_{1}=f\eta^{\epsilon}_{1}$ for some function $f$ and $g_{*}(R_{1}^{\epsilon})=R_{1}^{\epsilon}$ which implies that $f\equiv 1$. Consequently, $\eta^{\epsilon}_{1}$ induces a contact form $\eta$ on ${\widebar{M}}$. We claim that $\eta$ is the desired contact form satisfying (5.2).
As a matter of fact, note that the contractible periodic orbits of $\eta^{\epsilon}_{1}$ and $\eta$ are precisely the contractible iterations of the fibers over the fixed points of the torus action on $B$ (see [6, Proposition 3.9]) and that all these orbits are elliptic. Therefore, $\eta^{\epsilon}_{1}$ and $\eta$ are lacunary. It is easy to see that they are also non-degenerate. Moreover, we have that the map $\psi:{\mathscr{P}}^{0}(\eta^{\epsilon}_{1})\to{\mathscr{P}}^{0}(\eta)$ furnished by Lemma 5.1 is a bijection. Indeed, since the contractible periodic orbits of $\eta^{\epsilon}_{1}$ are precisely the contractible iterations of the fibers over the fixed points of the torus action on $B$, we have that the images of theses orbits coincide with the corresponding images of the orbits of the $T^{d+1}$-action and consequently the closed orbits of $\eta^{\epsilon}_{1}$ are symmetric with respect to the $G$-action. It easily follows from this that $\psi$ is injective. Therefore,
@@
${\mathrm{CC}}^{0}_{*}(\eta^{\epsilon}_{1})\cong{\mathrm{HC}}^{0}_{*}(M)\cong\bigoplus_{m\in\mathbb{N}}H_{\ast-2mc_{B}+n}(B;{\Lambda}),$ (5.3)
@@
Finally, let us prove that the lifts of the contractible closed orbits of $\alpha$ to $M$ are symmetric. Let $\{\gamma_{1},\dots,\gamma_{r_{B}}\}$ be the set of simple contractible orbits of $\alpha$ and $\{{\hat{\gamma}}_{1},\dots,{\hat{\gamma}}_{r_{B}}\}$ be lifts of these orbits, which are contractible closed orbits of the lifted contact form ${\hat{\alpha}}$ on $M$. These lifts are also simple contractible orbits. By Theorem 1.6, ${\hat{\alpha}}$ has precisely $r_{B}$ simple contractible closed orbits. If one of the orbits ${\hat{\gamma}}_{i}$ is not symmetric, we would have some $g\in G$ such that ${\text{Im}}({\hat{\gamma}}_{i})\neq{\text{Im}}(g({\hat{\gamma}}_{i}))$ and consequently we would have more than $r_{B}$ simple contractible closed orbits of ${\hat{\alpha}}$, a contradiction.
@@ 6. Multiplicity of periodic orbits for lens spaces and their unit cosphere bundles
As in Theorem 1.6 (see Remark 4.1), the only point in the proof of Theorem 1.1 in [27] that uses that $M$ is a prequantization is to achieve the isomorphism (2.1) with $c_{B}>n/2$ (when $B$ is spherically positive monotone). As proved in the previous section, this isomorphism holds for some prequantizations of orbifolds as in Theorem 1.13. In particular, as explained in the introduction, this isomorphism holds for lens spaces and their unit cosphere bundles. Therefore, we can conclude the following results.
@@
Note that in Theorem 6.1 we assume that $\alpha$ is index-admissible because a lens space does not admit, in general, a filling satisfying the assumption (F) of Theorem 1.6. This hypothesis is not necessary in Theorem 6.2 because $S^{*}L^{m}_{p}(\ell_{0},\dots,\ell_{d-1})$ satisfies this assumption. Moreover, note that the dimension of $S^{*}L^{m}_{p}(\ell_{0},\dots,\ell_{d-1})$ is $12m-1=2(m-1)+$ 1 with $m-1$ even since $m$ is odd.

### v3

@@ 4.2. Proof of the theorem
$\{1\leq i\leq r;\ {\mu}(\gamma_{i}^{k_{i}})<d-1\}+\#\{1\leq i\leq r;\ {\mu}(\gamma_{i}^{k_{i}})>d+1\}=r_{B}-\dim{\mathrm{H}}_{n}(B;\mathbb{Q}).$
@@
$\{1\leq i\leq r;\ {\mu}(\gamma_{i}^{k_{i}})<d-1\text{ or }{\mu}(\gamma_{i}^{k_{i}})>d+1\text{ or }{\mu}(\gamma_{i}^{k_{i}})=d\}=r.$
@@
Let us first prove that $\alpha$ has precisely $r_{B}$ contractible closed orbits. First of all, note that $\widebar{M}$} does not need to have a symplectic filling. Thus, we will argue as in the proof of Theorem 1.6 under the assumption (NF), but for this we have to show that $\alpha$ is index-admissible. Let $^{\hat{\alpha}}$ be the lift of $\alpha$ to $M$.
@@
[…]
@@
Let $^{\hat{\gamma}}$ be a contractible closed orbit of $^{\hat{\alpha}}$ and $\tau:M\to{\widebar{M}}$ be the quotient projection. Define $\psi({\hat{\gamma}})=\tau\circ{\hat{\gamma}}$. Clearly, this is a contractible closed orbit of $\alpha$ with the same index. To construct $\rho$, let $\gamma$ be a contractible closed orbit of $\alpha$ and choose a point $x_{0}\in\tau^{-1}(\gamma(0))$. Since $\tau:M\to{\widebar{M}}$ is a finite covering and $\gamma$ is contractible, we have that $\gamma$ admits a closed lift $^{\hat{\gamma}}$ such that ${\hat{\gamma}}(0)=x_{0}$. This lift is also contractible because given a capping disk $f$ of $\gamma$ (i.e. a continuous map $f:D^{2}\to{\widebar{M}}$ such that $f|_{\partial D^{2}}=\gamma$) we have that $f$ admits a lift $\hat{f}$ to $M$ such that $\hat{f}|_{\partial D^{2}}={\hat{\gamma}}$. Define $\rho(\gamma)={\hat{\gamma}}$. Then $\psi(\rho(\gamma))=\psi({\hat{\gamma}})=\gamma$ as desired. ∎
[…]
@@
for every contractible closed orbit $^{\hat{\gamma}}$ of $^{\hat{\alpha}}$. But, by (2.1),
@@
which implies that ${k_{\text{m}in}}=2c_{B}-n$. By hypothesis, $c_{B}\geq 2$ and therefore ${k_{\text{m}in}}>3-n$. Hence, $^{\hat{\alpha}}$ is index-admissible and so is $\alpha$, since the map $\psi$ in Lemma 5.1 is surjective.
@@
(5.1) ${\mathrm{HC}}^{0}_{\ast}({\widebar{M}})\cong\bigoplus_{m\in\mathbb{N}}{\mathrm{H}}_{\ast-2mc_{B}+n}(B;{\Lambda}).$
@@
(5.2) ${\mathrm{CC}}^{0}_{*}(\eta)\cong\bigoplus_{m\in\mathbb{N}}{\mathrm{H}}_{\ast-2mc_{B}+n}(B;{\Lambda})$
@@
Let ${\widebar{\xi}}=\tau_{*}\xi$ be the contact structure on $\widebar{M}$}, where $\tau:M\to{\widebar{M}}$, as before, is the quotient projection. To define the equivariant symplectic homology of contractible orbits ${\mathrm{HC}}^{0}_{*}({\widebar{M}})$ using the symplectization of $\widebar{M}$} (with an integral grading) it is enough to have that $c_{1}({\widebar{\xi}})|_{\pi_{2}({\widebar{M}})}=0$: we do not need that $c_{1}({\widebar{\xi}})|_{H_{2}({\widebar{M}};\mathbb{Q})}=0$ as in assumption (NF). Indeed, this assumption is needed in Theorem 1.6 in order to compute ${\mathrm{HC}}^{0}_{*}({\widebar{M}})$ in item (b) of Proposition 2.1. Here, we do not need this since we compute ${\mathrm{HC}}^{0}_{*}({\widebar{M}})$ directly from the construction of $\eta$. The condition $c_{1}({\widebar{\xi}})|_{\pi_{2}({\widebar{M}})}=0$ holds because, by the assumption that $(M,\xi)$ satisfies (F), we have that $c_{1}(\xi)|_{\pi_{2}(M)}=0$ which is equivalent to the condition $c_{1}({\widebar{\xi}})|_{\pi_{2}({\widebar{M}})}=0$ since the induced map $\tau_{\#}:\pi_{2}(M)\to\pi_{2}({\widebar{M}})$ is an isomorphism.
@@
Since $B$ is closed, this action has finitely many isotropy subgroups $G_{1},\dots,G_{m}$. At most one of these subgroups has dimension $d$, say $G_{m}$. Since the codimension of all other stabilizers $G_{1},\dots,G_{m-1}$ is positive, we can find a compact 1-parameter subgroup of $T^{d}$ which is not contained in any $G_{i}$ with $1\leq i<m$. The existence of this 1-parameter subgroup follows from the fact the union of the compact 1-parameter subgroups of $T^{d}$ is a dense subset of $T^{d}$. This is the desired subgroup. ∎
@@
where $R_{\beta}$ is the Reeb flow of $\beta$ (that generates the circle action of the prequantization $M$ whose orbits are the fibers) and $X_{i}^{h}$ is the horizontal lift of $X_{i}$; see [7, Lemma 3.4]. The flow of $R_{i}$ generates a circle action [7, Proposition 3.6] and commutes with the flow of $R_{\beta}$ [7, Lemma 3.8]. We also have that $R_{i},R_{j}]=0$ for every $i$ and $j$. As a matter of fact,
@@
where ${\hat{H}}_{i}R_{\beta},{\hat{H}}_{j}R_{\beta}]=0$ because ${\hat{H}}_{i}$ and ${\hat{H}}_{j}$ are invariant by the flow of $R_{\beta}$, ${\hat{H}}_{i}R_{\beta},X_{j}^{h}]=0$ (and similarly $X_{i}^{h},{\hat{H}}_{j}R_{\beta}]=0$) since $R_{\beta},X_{j}^{h}]=0$ (see [7, Lemma 3.8]) and $X_{j}^{h}({\hat{H}}_{i})=X_{j}(H_{i})=\{H_{j},H_{i}\}=0$, $X_{i},X_{j}]^{h}=0$ is the horizontal lift of $X_{i},X_{j}]$, and $\omega(X_{i},X_{j})=0$ because $X_{i}$ and $X_{j}$ commute.
@@
Now, let $\epsilon>0$ be an irrational number and define $\eta_{1}^{\epsilon}=\beta/({\hat{H}}_{1}+\epsilon)$ so that the Reeb vector field of $\eta_{1}^{\epsilon}$ is $R_{1}^{\epsilon}=R_{1}+\epsilon R_{d+1}$. By the previous discussion, we have that the $T^{d+1}$-action preserves $\eta^{\epsilon}_{1}$ as well: given $g\in T^{d+1}$ we have that $g^{*}\eta^{\epsilon}_{1}=f\eta^{\epsilon}_{1}$ for some function $f$ and $g_{*}(R_{1}^{\epsilon})=R_{1}^{\epsilon}$ which implies that $f\equiv 1$. Consequently, $\eta^{\epsilon}_{1}$ induces a contact form $\eta$ on $\widebar{M}$}. We claim that $\eta$ is the desired contact form satisfying (5.2).
As a matter of fact, note that the contractible periodic orbits of $\eta^{\epsilon}_{1}$ and $\eta$ are precisely the contractible iterations of the fibers over the fixed points of the torus action on $B$ (see [7, Proposition 3.9]) and that all these orbits are elliptic. Indeed, note that the contractible orbits of $\eta^{\epsilon}_{1}$ projects to contractible orbits of $\eta$ and, since $M\to{\widebar{M}}$ is a covering, every contractible orbit of $\eta$ is a projection a contractible orbit of $\eta^{\epsilon}_{1}$ (cf. Lemma 5.1); therefore, the contractible orbits of $\eta$ are elliptic if, and only if, so are the contractible orbits of $\eta^{\epsilon}_{1}$. But, since $R_{1}^{\epsilon}=({\hat{H}}_{1}+\epsilon)R_{\beta}+X_{1}^{h}$ and the vector fields $({\hat{H}}_{1}+\epsilon)R_{\beta}$ and $X_{1}^{h}$ commute, it is easy to see that every contractible orbit of $R_{1}^{\epsilon}$ is elliptic because every singularity of $X_{1}$ is elliptic. Therefore, $\eta^{\epsilon}_{1}$ and $\eta$ are lacunary. It is also easy to see that, choosing $\epsilon$ properly, they are also non-degenerate (because one can choose $\epsilon$ such that, for every fixed point $p$ of the $S^{1}$-action generated by $X_{1}$, all eigenvalues of $d\phi_{p}^{\epsilon}(p)$ are of the form $e^{2\pi i\theta}$ with $\theta\in\mathbb{R}\setminus\mathbb{Q}$, where $\phi_{p}^{\epsilon}$ is the time $1/(H_{1}(p)+\epsilon)$ map of the flow of $X_{1}$). Moreover, we have that the map $\psi:{\mathscr{P}}^{0}(\eta^{\epsilon}_{1})\to{\mathscr{P}}^{0}(\eta)$ furnished by Lemma 5.1 is a bijection. Indeed, since the contractible periodic orbits of $\eta^{\epsilon}_{1}$ are precisely the contractible iterations of the fibers over the fixed points of the torus action on $B$, we have that the images of theses orbits coincide with the corresponding images of the orbits of the $T^{d+1}$-action and consequently the closed orbits of $\eta^{\epsilon}_{1}$ are symmetric with respect to the $G$-action. It easily follows from this that $\psi$ is injective. Therefore,
@@
(5.3) ${\mathrm{CC}}^{0}_{*}(\eta^{\epsilon}_{1})\cong{\mathrm{HC}}^{0}_{*}(M)\cong\bigoplus_{m\in\mathbb{N}}H_{\ast-2mc_{B}+n}(B;{\Lambda}),$
@@
Finally, let us prove that the lifts of the contractible closed orbits of $\alpha$ to $M$ are symmetric. Let $\{\gamma_{1},\dots,\gamma_{r_{B}}\}$ be the set of simple contractible orbits of $\alpha$ and $\{{\hat{\gamma}}_{1},\dots,{\hat{\gamma}}_{r_{B}}\}$ be lifts of these orbits, which are contractible closed orbits of the lifted contact form $^{\hat{\alpha}}$ on $M$. These lifts are also simple contractible orbits. By Theorem 1.6, $^{\hat{\alpha}}$ has precisely $r_{B}$ simple contractible closed orbits. If one of the orbits ${\hat{\gamma}}_{i}$ is not symmetric, we would have some $g\in G$ such that ${\text{Im}}({\hat{\gamma}}_{i})\neq{\text{Im}}(g({\hat{\gamma}}_{i}))$ and consequently we would have more than $r_{B}$ simple contractible closed orbits of $^{\hat{\alpha}}$, a contradiction.
@@ 6. Multiplicity of periodic orbits for lens spaces and their unit cosphere bundles
As in Theorem 1.6 (see Remark 4.1), the only point in the proof of Theorem 1.1 in [30] that uses that $M$ is a prequantization is to achieve the isomorphism (2.1) (when $B$ is spherically positive monotone). As proved in the previous section, this isomorphism holds for some prequantizations of orbifolds as in Theorem 1.16. In particular, as explained in the introduction, this isomorphism holds for lens spaces and their unit cosphere bundles. Therefore, we can conclude the following results.
@@
Note that in Theorem 6.1 we assume that $\alpha$ is index-admissible because a lens space does not admit, in general, a filling satisfying the assumption (F) of Theorem 1.6. This hypothesis is not necessary in Theorem 6.2 because $S^{*}L^{m}_{p}(\ell_{0},\dots,\ell_{d-1})$ satisfies this assumption. Moreover, note that the dimension of $S^{*}L^{m}_{p}(\ell_{0},\dots,\ell_{d-1})$ is 2m1=2(m1)+12m-1=2(m-1)+1 with m1m-1 even since $m$ is odd.